OpenStudy (anonymous):

URGENT NEEDED A person buys a refrigerator for rupees 12000. He pays rupees 6000 in cash and agrees to pay the balance in 12 annual instalements of rupees 500 each. If the rate of interest is 12% and he pays it with instalements on the unpaid amount, how much did the refrigerator cost him ?

OpenStudy (anonymous):

does your question mean how much does it cost him to pay for the refrigerator or initially what is the cost

OpenStudy (anonymous):

\[\huge F=x\frac{(i+1)^n-1}{i}\]

OpenStudy (anonymous):

x=500 n=12*12=144 i=0.12/12=0.01

OpenStudy (anonymous):

\[P(1+i)^n=x \frac{(1+i)^n-1}{i}\] we can say this person took a loan of 12000-6000=6000 but how long did he pay this ????

OpenStudy (anonymous):

confused!!!

OpenStudy (kropot72):

The amount unpaid reduces by 500 rupees each year. The interest on 500 rupees for 1 year at 5% is 60 rupees. The interest for the first year is 6000 * 0.12 = 720 rupees. The interest payments form an arithmetic progression with the first term a = 720, the common difference d = -60 and the number of terms n = 12. The sum of the 12 terms S is given by \[S=\frac{n}{2}(2a+(n-1)d)=\frac{12}{2}(2\times 720+(12-1)-60)=4680\ rupees\] The cost of the refrigerator is the sum of : Initial payment = 6000 rupees twelve payments of 500 rupees = 6000 rupees Total interest payments = 4680 rupees

OpenStudy (kropot72):

@singhmmm Do you follow the method above?

OpenStudy (kropot72):

\[Total\ cost=6000+6000+4680=?\ rupees\]

OpenStudy (anonymous):

@kropot72 thank u so much

OpenStudy (kropot72):

You're welcome :)